# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ A.1 Proofs of the Auxiliary Lemmas and Resutls
$\|\bm{U}_{j,1}-\bm{U}_{j,2}\|^{2}_{\mathrm{op}}$
$\quad\leq\|\bm{C}_{j-1,1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j,1}\|^{2}_{\mathrm{op}}\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{1,1})^{-1}\|_{\mathrm{op}}\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}\bm{L}_{j,2})^{-1}\|_{\mathrm{op}}\|\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{j,1}$
$\qquad-\bm{L}^{\mathrm{T}}_{2,j}\bm{C}_{j-1,2}\bm{L}_{j,2}\|_{\mathrm{op}}+\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}bL_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{C}_{j-1,1}\bm{L}_{j,1}-\bm{C}_{j-1,2}\bm{L}_{j,2}\|^{2}_{\mathrm{op}}$
S3 Attributes of the realized graph
The nodal attributes are computed using NetworkToolbox [Christensen, 2018], and brainGraph [Watson, 2020] in R. For the convenience of the reader, we briefly describe the definitions and characteristics of the nodal attributes. For a graph $G$ with $V$ nodes, let us denote the adjacency matrix by $\bm{A}=(\!(a_{i,j})\!)_{1\leq i,j\leq V}$. Then, the degree of node $i$ is $\sum_{j}a_{i,j}$. The degree refers to the strength of each node in a graph. Global efficiency of the graph $G$ is measured by $F(G)=[V(V-1)]^{-1}\sum_{i\neq\ell\in G}r_{i,\ell}^{-1}$, where $r_{i,\ell}$ the shortest path length between the nodes $i$ and $\ell$. The nodal efficiency is defined by $N_{i}^{-1}\sum_{\ell\in N_{i}}F(G_{i})$, where $G_{i}$ is the subgraph of neighbors of $i$ and $N_{i}$ stands for the number of nodes in $G_{i}$. The corresponding betweenness centrality is $\sum_{\ell\neq i\neq k}{s_{\ell,k}(i)}/{s_{\ell,k}}$, where $s_{\ell,k}$ stands for the total number of the shortest paths from node $\ell$ to node $k$ and $s_{\ell,k}(i)$ denotes the number of those paths that pass through $i$. Nodal impact quantifies the impact of each node by measuring the change in average distances in the network upon removal of that node [Kenett et al., 2011]. The local shortest path length of each node is defined by the average shortest path length from that node to other nodes in the network. It is used to quantify the mean separation of a node in a network. Eccentricity is the maximal shortest path length between a node and any other node. Thus, it captures how far a node is from its most distant node in the network. Participant coefficient requires detecting communities first within the network. While using the R function participation of R package NetworkToolbox, we use the walktrap algorithm [Pons and Latapy, 2005] to detect communities. The participation coefficient quantifies the distribution of a node’s edges among different communities in the graph. It is zero if all the edges of a node are entirely restricted to its community, and it reaches its maximum value of 1 when the node’s edges are evenly distributed among all the communities. Within-community centrality is used to describe how central a node’s community is within the whole network. We use the walktrap algorithm to detect the communities. The final centrality measure that we include in our study is leverage centrality, introduced by Joyce et al. [2010]. Leverage centrality is the ratio of the degree of a node to its neighbors, and it thus measures the reliance of its immediate neighbors on that node for information.

### v2

@@ A.1 Proofs of the Auxiliary Lemmas
$\|\bm{U}_{j,1}-\bm{U}_{j,2}\|^{2}_{\mathrm{op}}$ $\leq\|\bm{C}_{j-1,1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j,1}\|^{2}_{\mathrm{op}}\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{1,1})^{-1}\|_{\mathrm{op}}$
×$\qquad\times\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}\bm{L}_{j,2})^{-1}\|_{\mathrm{op}}\|\bm{L}^{\mathrm{T}}_{j,1}\bm{C}_{j-1,1}\bm{L}_{j,1}-\bm{L}^{\mathrm{T}}_{2,j}\bm{C}_{j-1,2}\bm{L}_{j,2}\|_{\mathrm{op}}$
$\quad+\|(\bm{I}+\bm{L}^{\mathrm{T}}_{j,2}\bm{C}_{j-1,2}bL_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{C}_{j-1,1}\bm{L}_{j,1}-\bm{C}_{j-1,2}\bm{L}_{j,2}\|^{2}_{\mathrm{op}},$
establing the assertion. ∎
S3 Additional Figures
We present some additional figures from the data analysis.
Figure S4: Estimated change networks. The node numbers correspond to the variable names as described in McCracken and Ng (2020).